# Towards a theory of natural directed paths

### v3

@@ 2. Thick category of cubes
Let $\phi:[m]\to[n]$ be a map of $\mathcal{A}$. Let $\delta:[p]\to[m]\in\square$ such that $\phi\delta$ is one-to-one. Write $\phi\delta=\delta^{\prime}\phi^{\prime}$ with $\delta^{\prime}:[p]\to[n]\in\square$ and $\phi^{\prime}:[p]\to[p]\in\mathcal{A}$, the category of cubes $\mathcal{A}$ being thick by hypothesis. Since $\phi\delta$ is one-to-one, the set map $\phi^{\prime}:[p]\to[p]$ is one-to-one. By hypothesis, this implies that $\phi^{\prime}$ is the identity map. Thus $\phi\delta=\delta^{\prime}\in\square$. We have proved that $\phi\in\widehat{\square}$. ∎
To summarize, $\widehat{\square}$ is the greatest thick category of cubes for the inclusion not containing the symmetry maps and ^$\widehat{\square}_{S}$ is the greatest thick category of cubes for the inclusion.
@@
3. Metric and homotopical study
3.1 Notation.
@@ 3. Metric and homotopical study
Every morphism $f:[m]\to[n]$ of $\mathcal{A}$ is basic in the sense of [18, Definition 6.12] since every factorization of $f$ as a composite $[m]\to[p]\to[n]$ implies that $m\leqslant p\leqslant n$, and therefore that every factorization is not fundamental in the sense of [18, Definition 6.12]: $p<\min(m,n)=m$ is impossible indeed. Hence, for all $n\geqslant 0$, the subcategory $\mathcal{A}_{=n}$ of $\mathcal{A}$ generated by the objects of degree $n$ and the basic morphisms in the sense of [18, Definition 6.12] between them is the full subcategory of $\mathcal{A}$ having one object $[n]$. In particular, one has $\mathcal{A}_{=n}([n],[n])=\mathcal{A}([n],[n])$.
3.2 Proposition.
The small category $\mathcal{A}$ is c-Reedy in the sense of [18, Definition 8.25].
@@
Let $f$ be a map of $\mathcal{A}$. Consider the category of factorizations of $f$: its objects are the pairs of morphisms $(h,g)$ such that $hg=f$ with $h\in\overrightarrow{\mathcal{A}}$ and $g\in\overleftarrow{\mathcal{A}}$ and its morphisms $k:(h,g)\to(h^{\prime},g^{\prime})$ are morphisms $k$ (which are called connecting morphisms) such that there is a commutative diagram
[…]
By [11, Proposition 5.8] and since the factorization of Proposition 2.5 restricts to a factorization in $\mathcal{A}$, the category of factorization of $f$ has a final object. The rest of the proof goes like the proof of [11, Proposition 5.9]. Let us summarize the argument. One has $\overleftrightarrow{\mathcal{A}}\subset\overrightarrow{\mathcal{A}}\cap\overleftarrow{\mathcal{A}}$ (first axiom). Every morphism of $\overleftrightarrow{\mathcal{A}}$ is degree-preserving (second axiom). Every morphism of $\overrightarrow{\mathcal{A}}\backslash\overleftrightarrow{\mathcal{A}}$ strictly raises degree and every morphism of $\overleftarrow{\mathcal{A}}\backslash\overleftrightarrow{\mathcal{A}}=\varnothing$ strictly lowers degree (third axiom). The category of factorizations of $f$ with connecting maps in $\overleftrightarrow{\mathcal{A}}$ is connected since it has a final object (fourth axiom). For every $n\geqslant 0$, and any degree $m<n$, the functor $\overleftarrow{\mathcal{A}}([n],-):\mathcal{A}_{=m}\to{{\rm{\mathbf{Set}}}}$ is an (empty) coproduct of retracts of representables because $\mathcal{A}([n],[m])=\varnothing$ (fifth axiom). ∎
3.3 Notation.
@@
3.4 Notation.
Let $n\geqslant 0$. Following the notations of [18, page 37], let
@@
By Proposition 3.2 and [18, Theorem 8.26], the small category $\mathcal{A}$ is almost c-Reedy in the sense of [18, Definition 8.8]. The proof is complete thanks to [18, Theorem 8.9]. ∎
3.6 Proposition.
@@
$\partial_{n}\mathcal{A}([p],[q])=\begin{cases}\varnothing&\hbox{ if }p>q\hbox{or }n\leqslant p\\ \mathcal{A}([p],[q])&\hbox{ if }p\leqslant q\hbox{ and }p<n\end{cases}$
@@
It is mutatis mutandis the proof of [11, Proposition 5.13]: it suffices to change the category of cubes in the proof and to use Proposition 2.5 which restricts to a factorization in $\mathcal{A}$ by definition of a thick category of cubes. ∎
3.7 Theorem.
@@
The proof follows the road map of the proof of [11, Theorem 5.17] and makes use of Theorem 3.5 and Proposition 3.6. ∎
3.8 Notation.
The category of $\Delta$-generated spaces or of $\Delta$-Hausdorff $\Delta$-generated spaces (cf. [7, Section 2 and Appendix B]) is denoted by ${\mathbf{Top}}$
3.9 Definition.
[11, Definition 3.2] Let $f=(f_{1},\dots,f_{n}):[n]\to[n]$ be a cotransverse map. Let $\operatorname{T}(f):[0,1]^{n}\to[0,1]^{n}$ be the map defined by
@@
The three mappings $[n]\mapsto[0,1]^{n}$ for $n\geqslant 0$, $f:[n]\to[n]\in\mathcal{A}\mapsto\operatorname{T}(f)$ and $\delta_{i}^{\alpha}:[n-1]\to[n]\mapsto\operatorname{T}(\delta_{i}^{\alpha})$ for $n\geqslant 1$ give rise to a functor from $\mathcal{A}\subset\widehat{\square}_{S}$ to ${\mathbf{Top}}$ denoted by $|\mathcal{A}[*]|_{geom}$ by [11, Theorem 3.9] and to a functor from $\mathcal{A}\subset\widehat{\square}_{S}$ to $\mathbf{LvMet}$ denoted by $|\mathcal{A}[*]|_{\overrightarrow{d}\!_{1}}$ by [11, Theorem 3.16]. Let $K$ be an $\mathcal{A}$-set. Let
@@
These give rise to two colimit-preserving functors from $\mathcal{A}$-sets to topological spaces and Lawvere metric spaces respectively. The latter functor factoring as a composite $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}\to\widehat{\square}_{S}^{op}{{\rm{\mathbf{Set}}}}\to\mathbf{LvMet}$, we define the notion of (tame or not) natural $d$-path in the geometric realization $|K|_{geom}$ like in [11, Section 4]. In fact, by using the inclusion $\mathcal{A}\subset\widehat{\square}_{S}$ and the fact that $\mathcal{A}$ is thick, we can mimick all constructions of [11] and recover, thanks to Theorem 3.7, the results already proved for ^$\widehat{\square}_{S}$. More precisely, we obtain what follows.
3.11 Definition.
[3, Definition 4.11] A flow is a small semicategory enriched over the closed monoidal category $({\mathbf{Top}},\times)$. The corresponding category is denoted by ${{\rm{\mathbf{Flow}}}}$. The objects are called states and the morphisms execution paths.
There is an inclusion functor ${{\rm{\mathbf{PoSet}}}^{+}}\subset{{\rm{\mathbf{Flow}}}}$ such that there is an execution path from $\alpha$ to $\beta$ if and only if $\alpha<\beta$. Let $r\in\{q,m,h\}$. The category ${{\rm{\mathbf{Flow}}}}$ is equipped with its r-model structure [8, Theorem 7.4]. The weak equivalences of the r-model structure are the maps of flows $f:X\to Y$ inducing a bijection on states and a weak equivalence of the r-model structure of ${\mathbf{Top}}$ on execution paths.
3.12 Definition.
@@
The composite functor $\mathcal{A}\subset\widehat{\square}_{S}\to{{\rm{\mathbf{Flow}}}}$ taking $[n]$ to the flow $|\square_{S}[n]|_{nat}$ defined in [11, Proposition 7.1] induces a colimit-preserving functor
@@
from $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}$ to ${{\rm{\mathbf{Flow}}}}$ which is a m-realization functor by [11, Theorem 7.4]. The composite functor $\mathcal{A}\subset\widehat{\square}_{S}\to{{\rm{\mathbf{Flow}}}}$ taking $[n]$ to the flow ($(\{0<1\}^{n})^{cof}$, where $(-)^{cof}$ is a q-cofibrant replacement of ${{\rm{\mathbf{Flow}}}}$, induces a colimit-preserving functor
@@
from $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}$ to ${{\rm{\mathbf{Flow}}}}$ which is a q-realization functor by [11, Theorem 6.7].
3.14 Theorem.
@@
[…]
@@
We follow the proof of [11, Theorem 7.4]. Details are left to the reader. ∎
4. Ziemiański’s cube chain
@@
where the notation *** means that the final state $1_{n_{i}}$ of the $\mathcal{A}$-set $\mathcal{A}[n_{i}]$ is identified with the initial state $0_{n_{i+1}}$ of the $\mathcal{A}$-set $\mathcal{A}[n_{i+1}]$ for $11\leqslant i\leqslant p-$ 1.
4.3 Definition.
@@
${\mathcal{A}[\underline{n}_{a}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a}$${K\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}[\underline{n}_{b}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{b}$${K}$

### v4

@@ 2 Thick category of cubes
Let $\phi:[m]\to[n]$ be a map of $\mathcal{A}$. Let $\delta:[p]\to[m]\in\square$ such that $\phi\delta$ is one-to-one. Write $\phi\delta=\delta^{\prime}\phi^{\prime}$ with $\delta^{\prime}:[p]\to[n]\in\square$ and $\phi^{\prime}:[p]\to[p]\in\mathcal{A}$, the category of cubes $\mathcal{A}$ being thick by hypothesis. Since $\phi\delta$ is one-to-one, the function $\phi^{\prime}:[p]\to[p]$ is one-to-one. By hypothesis, this implies that $\phi^{\prime}$ is the identity map. Thus $\phi\delta=\delta^{\prime}\in\square$. We have proved that $\phi\in\widehat{\square}$. ∎
To summarize, $\widehat{\square}$ is the greatest thick category of cubes for the inclusion not containing the symmetry maps and $\widehat{\square}_{S}$ is the greatest thick category of cubes for the inclusion.
@@
3 Metric and homotopical study
For the ease of the reader, we recall a few definitions from c-Reedy before generalizing the results of DirectedDegeneracy . The point is not to overload this section but to make it comprehensible.
Definition 3.1.
(c-Reedy, , Definition 6.12) Let $\mathcal{C}$ be a category equipped with an ordinal degree function on its objects.
  • A morphism is level if its domain and codomain have the same degree.
  • The degree of a factorization $(h,g)$ of a morphism $f$ is the degree of the intermediate object (i.e. the domain of $h$ which is the codomain of $g$).
  • A factorization of a morphism $f$ is fundamental if its degree is strictly less than the degrees of both the domain and codomain of $f$.
  • A morphism is basic if it does not admit any fundamental factorization.
Definition 3.2.
(c-Reedy, , page 37) Let $\mathcal{C}$ be a category equipped with an ordinal degree function on its objects. The $\delta$-th stratum of $\mathcal{C}$, denoted by $\mathcal{C}_{=\delta}$, is the subcategory of $\mathcal{C}$ generated by the objects of degree $\delta$ and by the basic morphisms between them.
Definition 3.3.
Let $\mathcal{C}$ be a category. Let $f$ be a map of $\mathcal{C}$. The category of factorizations of $f$ has for objects the pairs $(h,g)$ such that $hg=f$ and for morphisms $k:(h,g)\to(h^{\prime},g^{\prime})$ the morphisms $k$ of $\mathcal{C}$ (which are called connecting morphisms) such that there is a commutative diagram
$\bullet$$\bullet$$\bullet$$\bullet$$\bullet$$\bullet$$\scriptstyle{g^{\prime}}$$\scriptstyle{h^{\prime}}$$\scriptstyle{g}$$\scriptstyle{k}$$\scriptstyle{h}$
Definition 3.4.
(c-Reedy, , Definition 8.25) A c-Reedy category $\mathcal{C}$ is a small category equipped with an ordinal degree function $d$ on its objects, and subcategories $\overleftrightarrow{\mathcal{C}}$, $\overrightarrow{\mathcal{C}}$ and $\overleftarrow{\mathcal{C}}$ containing all objects such that
  1. 1.
    $\overleftrightarrow{\mathcal{C}}\subseteq\overrightarrow{\mathcal{C}}\cap\overleftarrow{\mathcal{C}}$.
  2. 2.
    Every morphism in $\overleftrightarrow{\mathcal{C}}$ is level.
  3. 3.
    Every morphism in $\overrightarrow{\mathcal{C}}\backslash\overleftrightarrow{\mathcal{C}}$ strictly raises degree, and every morphism in $\overleftarrow{\mathcal{C}}\backslash\overleftrightarrow{\mathcal{C}}$ strictly lowers degree.
  4. 4.
    Every morphism $f$ factors as $\overrightarrow{f}\overleftarrow{f}$, where $\overrightarrow{f}\in\overrightarrow{\mathcal{C}}$ and $\overleftarrow{f}\in\overleftarrow{\mathcal{C}}$. The subcategory of the category of factorizations of $f$ generated by the pairs $(h,g)$ with $h\in\overrightarrow{\mathcal{C}}$ and $g\in\overleftarrow{\mathcal{C}}$ and such that the connecting morphisms belong to $\overleftrightarrow{\mathcal{C}}$ is connected for all $f$.
  5. 5.
    For any object $x$ and any degree $\delta<d(x)$, the functor $\overleftarrow{\mathcal{C}}(x,-):\overleftrightarrow{\mathcal{C}}_{=\delta}\to{{\rm{\mathbf{Set}}}}$ is a coproduct of retracts of representables.
Notation 3.5.
@@ 3 Metric and homotopical study
Every morphism $f:[m]\to[n]$ of $\mathcal{A}$ is basic since every factorization of $f$ as a composite $m]\to[p]\to[n]$ implies that $m\leqslant p\leqslant n$, and therefore that every factorization is not fundamental: $p<\min(m,n)=m$ is impossible indeed. Hence, for all $n\geqslant 0$, the $n$-th stratum $\mathcal{A}_{=n}$ is the full subcategory of $\mathcal{A}$ having one object $[n]$. Moreover, one has
$\mathcal{A}_{=n}([n],[n])=\mathcal{A}([n],[n])$
for all $n\geqslant 0$.
Proposition 3.6.
The small category $\mathcal{A}$ is c-Reedy.
@@
Let $f$ be a map of $\mathcal{A}$. Consider the subcategory $\mathcal{F}_{f}$ of the category of factorizations of $f$ generated by the pairs $(h,g)$ such that $h\in\overrightarrow{\mathcal{A}}$ and $g\in\overleftarrow{\mathcal{A}}$. Note that the connecting maps necessarily belong to $\overleftrightarrow{\mathcal{A}}$ since $\overleftrightarrow{\mathcal{A}}=\overleftarrow{\mathcal{A}}$. By (DirectedDegeneracy, , Proposition 5.8) and since the factorization of Proposition 2.5 restricts to a factorization in $\mathcal{A}$, the category $\mathcal{F}_{f}$ has a final object. The rest of the proof goes like the proof of (DirectedDegeneracy, , Proposition 5.9). Let us summarize the argument. One has $\overleftrightarrow{\mathcal{A}}\subset\overrightarrow{\mathcal{A}}\cap\overleftarrow{\mathcal{A}}$ (first axiom). Every morphism of $\overleftrightarrow{\mathcal{A}}$ is degree-preserving (second axiom). Every morphism of $\overrightarrow{\mathcal{A}}\backslash\overleftrightarrow{\mathcal{A}}$ strictly raises degree and every morphism of $\overleftarrow{\mathcal{A}}\backslash\overleftrightarrow{\mathcal{A}}=\varnothing$ strictly lowers degree (third axiom). The category $\mathcal{F}_{f}$ is connected since it has a final object (fourth axiom). For every $n\geqslant 0$, and any degree $m<n$, the functor $\overleftarrow{\mathcal{A}}([n],-):\mathcal{A}_{=m}\to{{\rm{\mathbf{Set}}}}$ is an (empty) coproduct of retracts of representables because $\mathcal{A}([n],[m])=\varnothing$ (fifth axiom). ∎
Notation 3.7.
@@
Notation 3.8.
Let $n\geqslant 0$. Following the notations of (c-Reedy, , page 37), let
@@
By Proposition 3.6 and (c-Reedy, , Theorem 8.26), the small category $\mathcal{A}$ is almost c-Reedy in the sense of (c-Reedy, , Definition 8.8). The proof is complete thanks to (c-Reedy, , Theorem 8.9). ∎
Proposition 3.10.
@@
[…]
@@
It is mutatis mutandis the proof of (DirectedDegeneracy, , Proposition 5.13): it suffices to change the category of cubes in the proof and to use Proposition 2.5 which restricts to a factorization in $\mathcal{A}$ by definition of a thick category of cubes. ∎
Theorem 3.11.
@@
The proof follows the road map of the proof of (DirectedDegeneracy, , Theorem 5.17) and makes use of Theorem 3.9 and Proposition 3.10. ∎
Notation 3.12.
The category of $\Delta$-generated spaces or of $\Delta$-Hausdorff $\Delta$-generated spaces (cf. (leftproperflow, , Section 2 and Appendix B)) is denoted by $\mathbf{Top}$}. There are three well-known model category structures on it, the $\{q,h,m\}$-model category structures.
Definition 3.13.
(DirectedDegeneracy, , Definition 3.2) Let $f=(f_{1},\dots,f_{n}):[n]\to[n]$ be a cotransverse map. Let $\operatorname{T}(f):[0,1]^{n}\to[0,1]^{n}$ be the map defined by
@@
The three mappings $n]\mapsto[0,1]^{n}$ for $n\geqslant 0$, $f:[n]\to[n]\in\mathcal{A}\mapsto\operatorname{T}(f)$ and $\delta_{i}^{\alpha}:[n-1]\to[n]\mapsto\operatorname{T}(\delta_{i}^{\alpha})$ for $n\geqslant 1$ give rise to a functor from $\mathcal{A}\subset\widehat{\square}_{S}$ to $\mathbf{Top}$} denoted by $|\mathcal{A}[*]|_{geom}$ by (DirectedDegeneracy, , Theorem 3.9) and to a functor from $\mathcal{A}\subset\widehat{\square}_{S}$ to $\mathbf{LvMet}$ denoted by $|\mathcal{A}[*]|_{\overrightarrow{d}\!_{1}}$ by (DirectedDegeneracy, , Theorem 3.16). Let $K$ be an $\mathcal{A}$-set. Let
@@
These give rise to two colimit-preserving functors from $\mathcal{A}$-sets to topological spaces and Lawvere metric spaces respectively. The latter functor factoring as a composite $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}\to\widehat{\square}_{S}^{op}{{\rm{\mathbf{Set}}}}\to\mathbf{LvMet}$, we define the notion of (tame or not) natural $d$-path in the geometric realization $|K|_{geom}$ like in (DirectedDegeneracy, , Section 4). In fact, by using the inclusion $\mathcal{A}\subset\widehat{\square}_{S}$ and the fact that $\mathcal{A}$ is thick, we can mimick all constructions of DirectedDegeneracy and recover, thanks to Theorem 3.11, the results already proved for $\widehat{\square}_{S}$. More precisely, we obtain what follows.
Definition 3.15.
(model3, , Definition 4.11) A flow is a small semicategory enriched over the closed monoidal category $({\mathbf{Top}},\times)$. The corresponding category is denoted by $\rm{\mathbf{Flow}}$}}. The objects are called states and the morphisms execution paths. The set of states of a flow $X$ is denoted by $X^{0}$. The space of execution paths from $\alpha$ to $\beta$ of a flow $X$ is denoted by $\mathbb{P}_{\alpha,\beta}X$.
There is an inclusion functor ${{\rm{\mathbf{PoSet}}}^{+}}\subset{{\rm{\mathbf{Flow}}}}$ such that there is an execution path from $\alpha$ to $\beta$ if and only if $\alpha<\beta$.
Let $r\in\{q,m,h\}$. The category $\rm{\mathbf{Flow}}$}} is equipped with its r-model structure (QHMmodel, , Theorem 7.4). The weak equivalences of the r-model structure are the maps of flows $f:X\to Y$ inducing a bijection on states and a weak equivalence of the r-model structure of $\mathbf{Top}$} for each map $\mathbb{P}_{\alpha,\beta}X\to\mathbb{P}_{f(\alpha),f(\beta)}Y$ for $(\alpha,\beta)$ running over $X^{0}\times X^{0}$. The fibrations of the r-model structure are the maps of flows $f:X\to Y$ inducing a fibration of the r-model structure of $\mathbf{Top}$} for each map $\mathbb{P}_{\alpha,\beta}X\to\mathbb{P}_{f(\alpha),f(\beta)}Y$ for $(\alpha,\beta)$ running over $X^{0}\times X^{0}$. The three model category structures of $\rm{\mathbf{Flow}}$}} are accessible in the sense of (zbMATH06722019, , Definition 5.1). The q-model structure of $\rm{\mathbf{Flow}}$}} is even combinatorial.
Definition 3.16.
@@
The composite functor $\mathcal{A}\subset\widehat{\square}_{S}\to{{\rm{\mathbf{Flow}}}}$ taking $[n]$ to the flow $|\square_{S}[n]|_{nat}$ defined in (DirectedDegeneracy, , Proposition 7.1) induces a colimit-preserving functor
@@
from $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}$ to $\rm{\mathbf{Flow}}$}} which is a m-realization functor by (DirectedDegeneracy, , Theorem 7.4). The composite functor $\mathcal{A}\subset\widehat{\square}_{S}\to{{\rm{\mathbf{Flow}}}}$ taking $[n]$ to the flow $(\{0<1\}^{n})^{cof}$, where $(-)^{cof}$ is a q-cofibrant replacement of $\rm{\mathbf{Flow}}$}}, induces a colimit-preserving functor
@@
from $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}$ to $\rm{\mathbf{Flow}}$}} which is a q-realization functor by (DirectedDegeneracy, , Theorem 6.7).
Theorem 3.18.
@@
$\mathbb{P}_{\alpha,\beta}|K|_{q}$$\mathbb{P}_{\alpha,\beta}F(K)$$\mathbb{P}_{\alpha,\beta}|K|_{nat}.$$\scriptstyle{\simeq}\scriptstyle{\simeq}$
@@
We follow the proof of (DirectedDegeneracy, , Theorem 7.4). Details are left to the reader. ∎
4 Ziemiański’s cube chain
@@
where the notation * means that the final state $1_{n_{i}}$ of the $\mathcal{A}$-set $\mathcal{A}[n_{i}]$ is identified with the initial state $0_{n_{i+1}}$ of the $\mathcal{A}$-set $\mathcal{A}[n_{i+1}]$ for $1\leqslant i\leqslant p-1$.
The category of posets being cocomplete, being locally presentable, the vertices of the $\underline{n}$-cube can be equipped with the poset structure induced by the one on each $\mathcal{A}[n_{i}]_{0}=\{0<1\}^{n_{i}}$.
The set $\operatorname{Vert}(\underline{n})$ is in bijection with the set consisting of the initial states of the $\mathcal{A}[n_{i}]$ for $1\leqslant i\leqslant p$ and the final state of $\mathcal{A}[n_{p}]$ viewed as subobjects of $\mathcal{A}[\underline{n}]$. For all $\in\operatorname{Vert}(\underline{n}), pqp\leqslant q$ as integers if and only if $p\leqslant q$ for the poset structure on the vertices of $\mathcal{A}[\underline{n}]$.
Using the fact that the $\underline{n}$-cube is a colimit, and $\mathbf{LvMet}$ being cocomplete, the set of vertices of the $\underline{n}$-cube can also be equipped with the $\overrightarrow{d}\!_{1}$ metric. Proposition 4.3 summarizes what is necessary to know about this Lawvere metric for the proof of Theorem 4.6.
Proposition 4.3.
Let $\underline{n}\in\operatorname{Seq}(n)$ with $n\geqslant 0$ as above. The following properties hold:
  1. 1.
    Let $x$ and $y$ be two vertices of $\mathcal{A}[\underline{n}]$ belonging to the same $\mathcal{A}[n_{i}]$. Then $\overrightarrow{d}\!_{1}(x,y)$ is the distance in $\mathcal{A}[n_{i}]$ from $x$ to $y$.
  2. 2.
    Let $p,q\in\operatorname{Vert}(\underline{n})$ which are identified with the corresponding vertices of $\mathcal{A}[\underline{n}]$ as explained above. Then $p\leqslant q$ implies $\overrightarrow{d}\!_{1}(p,q)=q-p$ and $p>q$ implies $\overrightarrow{d}\!_{1}(p,q)=\infty$.
  3. 3.
    If $x\in\mathcal{A}[n_{i}]_{0}$ and $y\in\mathcal{A}[n_{j}]_{0}$ with $i<j$, then
    $\overrightarrow{d}\!_{1}(x,y)=\overrightarrow{d}\!_{1}(x,n_{1}+\dots+n_{i})+n_{j-1}-n_{i}+\overrightarrow{d}\!_{1}(n_{1}+\dots+n_{j-1},y)$
    where $n_{1}+\dots+n_{i}\in\operatorname{Vert}(\underline{n})$ is identified with the final state of $\mathcal{A}[n_{i}]$ in $\mathcal{A}[\underline{n}]$ and where $n_{1}+\dots+n_{j-1}\in\operatorname{Vert}(\underline{n})$ is identified with the initial state of $\mathcal{A}[n_{j}]$ in $\mathcal{A}[\underline{n}]$.
Proof.
The category $\mathbf{LvMet}$ being the category of $([0,\infty],\geqslant,+,0)$-enriched small categories, we obtain for two vertices $x$ and $y$ of $\mathcal{A}[\underline{n}]$ the general formula
$\overrightarrow{d}\!_{1}(x,y)=\min_{\begin{subarray}{c}r\geqslant 1\\ (k_{1}, \dots,k_{r})\in\{1,\dots,p\}^{r}\end{subarray}}\min_{\begin{subarray}{c}(x_{0},\dots,x_{r})\\ x=x_{0},y=x_{r}\\ \forall i,\{x_{i},x_{i+1}\}\subset\mathcal{A}[n_{k_{i}}]_{0}\end{subarray}}\overrightarrow{d}\!_{1}(x_{0},x_{1})+\dots+\overrightarrow{d}\!_{1}(x_{r-1},x_{r}).$
Let $x$ and $y$ be two vertices of $\mathcal{A}[\underline{n}]$ belonging to the same $\mathcal{A}[n_{i}]$. Then $\overrightarrow{d}\!_{1}(x,y)$ in $\mathcal{A}[n_{i}]$ is the minimum by the triangular inequality. Hence $(1)$. Let $p,q\in\operatorname{Vert}(\underline{n})$ with $p\leqslant q$. For an $s$-cube $\mathcal{A}[s]$, $\overrightarrow{d}\!_{1}(0_{s},1_{s})=s$. We deduce that for any $r\geqslant 1$, any $(k_{1}, \dots,k_{r})\in\{1,\dots,p\}^{r}$ and any $(p=x_{0},\dots,x_{r}=q)$ such that $\forall i,\{x_{i},x_{i+1}\}\subset\mathcal{A}[n_{k_{i}}]_{0}$, the sum $\overrightarrow{d}\!_{1}(x_{0},x_{1})+\dots+\overrightarrow{d}\!_{1}(x_{r-1},x_{r})$ is equal to qpq-p or to $\infty$. Hence the first part of $(2)$. Assume now that $q>p$. Then in any sum of the type $\overrightarrow{d}\!_{1}(x_{0},x_{1})+\dots+\overrightarrow{d}\!_{1}(x_{r-1},x_{r})$ above, one of the terms is equal to $\infty$. Hence $(2)$. We deduce $(3)$ thanks to the following calculation:
$\overrightarrow{d}\!_{1}(x,y)$
$=\overrightarrow{d}\!_{1}(x,n_{1}+\dots+n_{i})+\overrightarrow{d}\!_{1}(n_{1}+\dots+n_{i},n_{1}+\dots+n_{j-1})+\overrightarrow{d}\!_{1}(n_{1}+\dots+n_{j-1},y)$
$=\overrightarrow{d}\!_{1}(x,n_{1}+\dots+n_{i})+n_{j-1}-n_{i}+\overrightarrow{d}\!_{1}(n_{1}+\dots+n_{j-1},y).$
Definition 4.4.
@@
$\mathcal{A}[\underline{n}_{a}]$${\mathcal{A}[\underline{n}_{b}]}K{K}$$\scriptstyle{a}$$\scriptstyle{b}$